OpenStudy (anonymous):

a contractor wants to use 34 feet of molding, cut into three pieces, to trim the sides and top of a garage door. The piece is 1.5 feet longer than three times the length of each shorter piece. Find the length of each piece?

OpenStudy (anonymous):

l=long piece s=short piece l+2s=34 1.5+3s=l

OpenStudy (anonymous):

Its a system of equations, do you know how to solve?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

Solve for l in one equation. l+2s=34 l=-2s+34 since l=-2s+34 put that in the other equation where l is 1.5+3s=-2s+34 now solve for s

OpenStudy (anonymous):

does s = 6.5?

OpenStudy (anonymous):

Yep, now you know s so plug that into one of the original equations and solve for l. l+2(6.5)=34

OpenStudy (anonymous):

so, l = 21?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

thanks!!